OpenStudy (anonymous):

Help!! A mixing blade on a food processor extends out 3 inches from its center. If the tip of the blade is turning at 600 rpm, what is the linear velocity of the tip of the blade in feet per minute?

OpenStudy (sidsiddhartha):

3 inch=.25 feet So the distance/rev = 2πr = 2π(.25) = π/2 feet v=ft per rev*rev per min=300*pi per min

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=20100715005549AA9EhzN lol copy cat

OpenStudy (sidsiddhartha):

yeah i've done it before- the problem :)